OpenStudy (jenniferfournierr):

solve x+2y=13 4x+8y=-9 choices (3,5) (-1,7) no solutions infinitely many solutions

MsBrains (ms-brains):

Any Ideas? :)

OpenStudy (jenniferfournierr):

no im really lost

MsBrains (ms-brains):

Steps: I will try to solve your system of equations. x+2y=13;4x+8y=−9 Step: Solvex+2y=13for x: x+2y+−2y=13+−2y(Add -2y to both sides) x=−2y+13 Step: Substitute−2y+13forxin4x+8y=−9: 4x+8y=−9 4(−2y+13)+8y=−9 52=−9(Simplify both sides of the equation) 52+−52=−9+−52(Add -52 to both sides) 0=−61 Answer: No solutions. _ Source https://www.mathpapa.com/algcalc2/home/?utm_expid=69051716-6.REXekKgVRPOx9_r9ROC2DA.2&utm_referrer=https%3A%2F%2Fwww.google.com%2F

OpenStudy (jenniferfournierr):

thank you so much can u help with another?

OpenStudy (jenniferfournierr):

solve 3x-5y=17 9x-15y=-4

TheSmartOne (thesmartone):

Multiply the first equation on both sides by 3. What do you get?

OpenStudy (jenniferfournierr):

9x-15y=17

MsBrains (ms-brains):

Clearly, tso will be helping you with this. So I'm out. :P

OpenStudy (jenniferfournierr):

thank u for your help !

TheSmartOne (thesmartone):

You have to multiply 3 on both sides. You fogot to multiply 17 by 3. You would have gotten 9x - 15 = 51 The two equations right next to each other are 9x - 15y = 51 9x - 15y = -4 What would happen if you subtract the second equation from the first one? What do you get? :)

OpenStudy (jenniferfournierr):

im not sure

TheSmartOne (thesmartone):

For example 2x - 5y = 2 3x + 5y = 3 If we subtracted, we would get: 2x - 5y - (3x + 5y) = 2 - 3 And then you simplify it It's basically like (a much more easier example) x = 5 y = 4 subtract both equations x - y = 5 - 4 and when you simplify it x - y = 1 Which makes sense if x = 5 and y = 4

OpenStudy (jenniferfournierr):

the choices are (-1,-4) (4,-1) no solutions infinitely many solutions

OpenStudy (jenniferfournierr):

would it be d then??

TheSmartOne (thesmartone):

Nope! Remember if you get something like 3 = 3, (which is always true), then you have infinitely many solutions. If you get something like 1 = 3, (which is false), then you have no solutions If you get something like x = 1, or y = 3, then you have one solution You can NEVER have two solutions in a system of linear equations

OpenStudy (princeharryyy):

The lines are parallel. hence no solution exists.

OpenStudy (princeharryyy):

For any lines to be parallel with no solutions the slopes of both the lines should be equal and the constant should be different.

TheSmartOne (thesmartone):

*facepalm*

OpenStudy (princeharryyy):

For the lines to have infinitely many solutions. 1. lines should be parallel i.e. same slope 2. the constant of both the equations should be same.

OpenStudy (princeharryyy):

Go with C no solutions. :)

TheSmartOne (thesmartone):

I fail to see how giving the asker the answer helps him learn how to do it.

OpenStudy (princeharryyy):

@TheSmartOne I think I gave u the explanation. Perhaps u don't have to solve every problem to find the answer. It's easily visible when u look at slopes and constans of the equations.

OpenStudy (jenniferfournierr):

im a girl and tso u didnt even get it right so

OpenStudy (princeharryyy):

constants*

OpenStudy (jenniferfournierr):

and it does help so i can look at what they told me and see the steps to get the right answer THNAKKSSS PRINCE HARRY

OpenStudy (princeharryyy):

well, pretty good with me. hv a good luck solving problems u need not solve for answers :)

TheSmartOne (thesmartone):

Yes, I was able to see that it had no solutions right away BUT the asker will not be able to see it without practice. That's why I asked her to multiply the equations by 3 and to subtract the equations. Both variables would be eliminated and an untrue equation would remain which would show that there was no solution. The first question, she got an answer. I highly doubt she understood what was happening. She took the answer and asked for help on the next question.